# Community Discovery Algorithm Based on Spatio-temporal Graph Embedding in Dynamic Social Networks

### v1

@@ Abstract
Community discovery is one of the key issues in the study of dynamic social networks. Traditional community discovery algorithms only focus on the establishment and disconnection of connections between nodes, failing to capture deeper factors. To address this limitation, in this work, we propose a community discovery algorithm based on spatiotemporal graph embedding (CDA-SGE), which integrates spatial information and evolutions of nodes to comprehensively capture the dynamic features of networks. Specifically, this algorithm employs Graph Convolutional Neural Networks (GCN) to aggregate latent spatial information, effectively representing the embedding of nodes in space. Temporal evolutions of the nodes are then modeled using Gated Recurrent Units (GRU), thereby solving problems such as node dynamism and relationship transmission. Finally, a Self-Organizing Map (SOM) is applied to cluster dynamic network representations and identify community affiliations of nodes. We then perform simulations on four types of dynamic networks and show that the CDA-SGE outperforms traditional community discovery algorithms in terms of purity, standardized mutual information, heterogeneity, and homogeneity. These results demonstrate the algorithm’s superior ability to accurately uncover community structures hidden in dynamic social networks.
@@
Community Discovery; Dynamic Social Networks; Spatiotemporal Graph Embedding; Gate Controlled Cycle Unit
journal: Graphical Models
[…]
[1]organization=Southern Methodist University, addressline=6425 Boaz Lane, postcode=75205, city=Dallas, TX, country=USA
@@ 1 Introduction
Community discovery(Kang et al., 2024), one of the key tasks in social network analysis, aims to identify community structures in the network that exhibit strong internal connections while being relatively independent of other groups. The applications of community discovery span a wide range of domains, providing valuable insights and effective solutions. In targeted marketing and personalization, communities offer a deeper understanding of shared interests and behavioral patterns(Paliouras, 2012), enabling the development of more effective recommendation systems and customized advertising strategies. In epidemiology and public health(Leung et al., 2004; Rostami et al., 2023), identified communities can represent populations at higher risk or groups vulnerable to the spread of infectious diseases, facilitating the design of targeted intervention and containment measures. Similarly, in security and fraud detection, community discovery plays a crucial role in identifying anomalous groups within financial or communication networks(Sarvari et al., 2014; Andre et al., 2012; Waskiewicz, 2012), helping analysts detect fraudulent activities, cyber threats, or other malicious entities. Finally, in social science and behavioral analysis(Dhawan et al., 2021; Lancieri and Durand, 2006), understanding community dynamics provides critical insights into cultural trends, patterns of social influence, and the evolution of social norms over time. These diverse applications underscore the importance of studying communities within dynamic networks to address real-world challenges.
In static social networks, numerous community discovery algorithms(Zhang et al., 2024) have achieved certain results. However, the temporal variability of dynamic social networks poses significant challenges for community discovery(He et al., 2024; Wu et al., 2024). Traditional static algorithms(Zhang et al., 2020) face challenges in capturing dynamic changes in network properties across different time points, and fail to accurately reflect the real community structure and its evolution process. Recently, graph embedding techniques(Yuan et al., 2022) emerged in the field of social network analysis. Such techniques map nodes in the network to a low-dimensional vector space, thereby preserving the structure and semantic information of nodes. However, most existing graph embedding methods still focus primarily on static networks, with significant limitations in the joint processing of temporal and spatial dimension information in dynamic social networks(Duan et al., 2021).
@@
This paper is organized as follows. In Sect. 1, we briefly discuss the motivation to propose a dynamic community discovery algorithm. In Sect. 2, we review some of the progress and limitations of current community discovery algorithms. In Sect. 3, we discuss the necessary definitions related to CDA-SGE and introduce the algorithm. Experiments and comparisons are conducted in Sect. 4. And finally Sect. 5 concludes the paper.
@@ 2 Related Works
Researchers have made significant progress in the study of community discoveries within dynamic social networks. Yang and Duan (2023) proposed a dynamic community discovery algorithm based on spectral clustering (SC-DCDA) to address the problems of network structure analysis, insufficient clustering information, and excessive computational overhead caused by the lack of historical information necessary for community nodes in static community discovery algorithms for social networks. This algorithm is based on the division of community network evolution events and the analysis of major community events. Traditional community discovery methods assume that nodes have an equal status in the network, ignoring the influence of nodes in community formation. To address this issue, Yang et al. (2023) proposed a community discovery method based on node influence expansion. This approach begins by calculating the local influence of the nodes using Monte Carlo simulation. Next, a novel eccentricity-based method is introduced to screen bridge nodes and enhance seed quality. Finally, dynamic programming is applied to expand the seeds and generate communities. To reveal hidden dynamic phenomena in virtual social networks, Li et al. (2021) proposed a multi-time dense subgraph discovery algorithm for online social network discovery, which considers both topic and time. In this algorithm, conversations or comments are modeled as an interactive network, which is then partitioned into topic-specific communities using its topological structure. Within each community, dense subgraphs are mined on the basis of time dimension, capturing temporal dynamics.
@@
However, current dynamic social network community discovery algorithms still have many limitations: (a) The explosive growth of social network data increases computational complexity, memory usage, and runtime. Many algorithms require recalculating at each time step, causing the time complexity to rise significantly with the number of steps, hindering real-time performance. (b) Some algorithms neglect temporal correlations and fail to capture the true evolution of communities and complex structures such as nesting, overlap, and hierarchy, leading to the loss of critical information. (c) Algorithms often struggle with the dynamic nature of social networks, exhibiting poor adaptability to structural changes and sensitivity to parameter adjustments. In addition, they are vulnerable to noise and abnormal data, compromising accuracy and reliability. Considering these limitations, to fully extract the dynamic features of the network and the interaction information between nodes, and to deeply mine the dynamic features of the network, we propose a community discovery algorithm based on spatiotemporal graph embedding (CDA-SGE). This algorithm employs Graph Convolutional Neural Networks (GCN) to aggregate latent spatial information, effectively representing the embedding of nodes in space. The temporal evolutions of nodes are then modeled using Gated Recurrent Units (GRU), thereby solving problems such as node dynamism and relationship transmission. Finally, a Self-Organizing Map (SOM) is applied to cluster dynamic network representations and identify community affiliations of nodes.
@@ 3.1 Definitions
Definition 1.
(Dynamic Network) Let $V=\{v_{1},v_{2},\cdots,v_{n}\}$ be the set of nodes in a dynamic network, and $E=\left\{e_{ij}\mid\forall i,j\in\{1,\ldots,n\},i\neq j\right\}$ be the set of edges. Then $g_{t}=(V,E_{t})$ is the snapshot network at time $t$. Here, the number of nodes remains constant, and the number of edges changes over time. A dynamic network $G$ consists of time snapshots of a network, i.e. $G=(V,E)=\left\{g_{1},g_{2},\ldots,g_{t},\ldots,g_{T}\right\},1\leq t\leq T$.
Definition 2.
(Dynamic Community) We use $C_{t}=\left\{c_{t}^{1},c_{t}^{2},\ldots,c_{t}^{p}\mid\forall t\in\{1,\ldots,T\}\right\}$ to denote the $p$ communities divided by the time slice network $g_{t}$. The dynamic community divided by the dynamic network $G$ is represented by $C=\left\{C_{1},C_{2},\ldots,C_{T}\right\}$.
Definition 3.
(Graph Embedding) Graph embedding can map dynamic network $G(V,E)$ from high-dimensional space to low-dimensional space. The mapping function $\Gamma$ at time $t$ is defined as
$\Gamma:v_{t}^{i}\longrightarrow f_{t}^{i}\in\mathbb{R}^{d},1\leq i\leq n,1\leq t\leq T,d\ll n.$
Here, $f_{t}^{i}$ is the embedding vector of $\Gamma$ to node $v_{i}$ in the network at time $t$ , and $d$ represents the dimension of the embedding vector, which is much smaller than the total number of nodes $n$.
Tab. 1 provides commonly used symbol representations and their corresponding descriptions in this paper.
@@
$G$ Dynamic network (dynamic graph)
$t$ current time
$T$ The number of time slice networks
@@
$n$ Number of nodes
$d$ Embedding vector dimension
$C$ Community Collection of Dynamic
$F_{S}$ Networks
$F_{t}$ Time domain embedding features at time $t$
Table 1: Symbols and descriptions used in this paper.
3.2 The CDA-SGE Algorithm
The dynamic network community discovery process based on spatiotemporal graph embedding, shown in Fig. 1, consists of three main components: spatial feature extraction, temporal feature extraction, and community discovery. In order to fully utilize the spatial information of the nodes in the network, we first employ GCN to learn the feature representation of the nodes from a spatial perspective and aggregate the social information of users who communicate with each other. Next, we share the GCN’s weight parameters using GRU to maintain the dynamics of spatial nodes and extract the features of dynamic nodes, addressing node dynamics and long-term dependencies to form the feature representation of the dynamic network. Finally, the SOM algorithm is applied to analyze the network embedding and perform community discovery for dynamic artificial and real social networks. We dub this process as the CDA-SGE algorithm.
Figure 1: Research framework of CDA-SGE.
@@ 3.3 Spatial Feature Extraction
In the spatial feature extraction stage, we start with node interaction learning and use GCN to learn the topological structure information of nodes in the snapshot network and obtain the embeddings of nodes. This process involves two key aspects: space embedding learning and weight parameter sharing.
@@ 3.3.1 Space Embedding Learning
GCN is composed of multiple layers of graph convolutions, which utilize convolution operations for neighborhood aggregation. At time $t$, in the $l$-th layer, the adjacency matrix $A_{t}$ and the embedding matrix $F_{t}^{l-1}$ are taken as input. The weight matrix $W_{t}^{l-1}$ is used to update the node embedding matrix to $F_{t}^{l}$, which is the output:
$F_{t}^{l}=\Gamma\left(F_{t}^{l-1},A_{t},W_{t}^{l-1}\right)=\sigma\left(\hat{A}_{t}F_{t}^{l-1}W_{t}^{l-1}\right).$ (1)
Here, $\hat{A}_{t}$, the normalized adjacency matrix, serves as an approximate graph convolution filter. It can be obtained by
$\hat{A}_{t}=\hat{D}_{t}^{-1/2}\hat{A}_{t}\hat{D}_{t}^{-1/2},$ (2)
$\hat{A}_{t}=A_{t}+\mathbb{I},$
where $\hat{A}_{t}$ is the adjacency matrix of the time snapshot network at time $t$ and $A=\left\{A_{1},A_{2},\ldots,A_{t},\ldots,A_{T}\right\}$ is the adjacency matrix composed of snapshot networks. If there is a connecting edge between node $v_{i}$ and $v_{j}$, $A_{ij}=0$ and $1$ otherwise. $\mathbb{I}$ is the identity matrix, which is added and summed with the adjacency matrix to form a self-loop among the nodes in the snapshot network, providing a more detailed representation of the characteristics of the nodes. $\hat{D}_{t}$ is the degree matrix representing the nodes of the network,
$\hat{D}_{t}=\operatorname{diag}\left(\sum_{j=1}^{n}A_{t_{(ij)}}\right).$ (3)
$F_{t}=\left\{f_{t}^{1},f_{t}^{2},\ldots,f_{t}^{n}\right\},1\leq t\leq T$ represents the characteristics of the node. In this paper, since the relevant dynamic network only includes nodes and edges, we use a random initialization method to generate input features from the nodes. The activation function $\sigma$ used in this paper follows LeakyReLU method.
In Eq. 2, at time $t$, the input of the first layer of the network is a randomly initialized node embedding matrix $F_{0}^{t}$, where one row represents the $d$-dimensional eigenvector of a node. After convolution in the $l$-th layer, the output features of each time snapshot aggregate the interaction information between the node and its neighbors.
@@ 3.3.2 Weight Parameter Sharing
Given the dynamic nature of the network, dynamic update strategies are incorporated into the graph convolutional layers. This is necessary because any change in nodes or edges triggers an adjustment of the weight parameters in the convolutional layers to adapt to the evolving network structure. Therefore, we follow the method of EvolveGCN(Duan et al., 2021) and use GRU to update the weight parameters of GCN at the previous time step, which serves as the weight of GCN at the current time step, to adapt to the development of dynamic networks. The weight update strategy for the $l$-th layer at time t is as follows:
@@
The GCN module aggregates information between nodes and their neighbors, while the GRU adaptively adjusts weight parameters over time. Together, these modules dynamically capture the evolving neighborhood information of the nodes. Thus, operations in the spatial feature extraction stage can be formally represented as
@@
Figure 2: Spatial feature extraction.

### v2

# GTENN: A Spatiotemporal Graph Neural Framework for Community Discovery in Dynamic Social Networks

@@ Abstract
Community discovery is one of the key issues in the study of dynamic social networks. Traditional community discovery algorithms mainly focus on the formation and dissolution of links between nodes, and thus fail to capture richer spatial and temporal patterns underlying network evolution. To address this limitation, we propose GTENN, a spatiotemporal graph neural framework for community discovery in dynamic social networks. GTENN integrates spatial structure and temporal dynamics within a unified embedding architecture. First, Graph Convolutional Networks (GCN) are employed to aggregate latent spatial information and learn expressive node representations at each snapshot. Next, Gated Recurrent Units (GRU) are used to model temporal evolutions of node embeddings, effectively capturing node dynamism and relationship propagation across time. Finally, a Self-Organizing Map (SOM) is applied to the learned spatiotemporal embeddings to cluster nodes and infer their community affiliations. We conduct experiments on four types of dynamic networks, and the results show that GTENN consistently outperforms traditional community discovery algorithms in terms of purity, normalized mutual information, homogeneity, and completeness. These findings demonstrate the superior ability of GTENN to accurately uncover evolving community structures hidden in dynamic social networks.
@@
Community Discovery; Dynamic Social Networks; Spatiotemporal Graph Embedding; Recurrent Unit; Self-organizing Map
@@ 1 Introduction
Community discovery[1], one of the key tasks in social network analysis, aims to identify community structures in the network that exhibit strong internal connections while being relatively independent of other groups. The applications of community discovery span a wide range of domains, providing valuable insights and effective solutions. In targeted marketing and personalization, communities offer a deeper understanding of shared interests and behavioral patterns[2], enabling the development of more effective recommendation systems and customized advertising strategies. In epidemiology and public health[3, 4], identified communities can represent populations at higher risk or groups vulnerable to the spread of infectious diseases, facilitating the design of targeted intervention and containment measures. Similarly, in security and fraud detection, community discovery plays a crucial role in identifying anomalous groups within financial or communication networks[5, 6, 7], helping analysts detect fraudulent activities, cyber threats, or other malicious entities. Finally, in social science and behavioral analysis[8, 9], understanding community dynamics provides critical insights into cultural trends, patterns of social influence, and the evolution of social norms over time. These diverse applications underscore the importance of studying communities within dynamic networks to address real-world challenges.
In static social networks, numerous community discovery algorithms[10] have achieved certain results. However, the temporal variability of dynamic social networks poses significant challenges for community discovery[11, 12]. Traditional static algorithms[13] face challenges in capturing dynamic changes in network properties across different time points, and fail to accurately reflect the real community structure and its evolution process. Recently, graph embedding techniques[14] emerged in the field of social network analysis. Such techniques map nodes in the network to a low-dimensional vector space, thereby preserving the structure and semantic information of nodes. However, most existing graph embedding methods still focus primarily on static networks, with significant limitations in the joint processing of temporal and spatial dimension information in dynamic social networks[15].
@@
This paper is organized as follows. In Sect. 1, we briefly discuss the motivation to propose a dynamic community discovery algorithm. In Sect. 2, we review some of the progress and limitations of current community discovery algorithms. In Sect. 3, we discuss the necessary definitions related to GTENN and introduce the algorithm. Experiments and comparisons are conducted in Sect. 4. And finally Sect. 5 concludes the paper.
@@ 2 Related Works
Researchers have made significant progress in the study of community discoveries within dynamic social networks. [16] proposed a dynamic community discovery algorithm based on spectral clustering (SC-DCDA) to address the problems of network structure analysis, insufficient clustering information, and excessive computational overhead caused by the lack of historical information necessary for community nodes in static community discovery algorithms for social networks. This algorithm is based on the division of community network evolution events and the analysis of major community events. Traditional community discovery methods assume that nodes have an equal status in the network, ignoring the influence of nodes in community formation. To address this issue, [17] proposed a community discovery method based on node influence expansion. This approach begins by calculating the local influence of the nodes using Monte Carlo simulation. Next, a novel eccentricity-based method is introduced to screen bridge nodes and enhance seed quality. Finally, dynamic programming is applied to expand the seeds and generate communities. To reveal hidden dynamic phenomena in virtual social networks, [18] proposed a multi-time dense subgraph discovery algorithm for online social network discovery, which considers both topic and time. In this algorithm, conversations or comments are modeled as an interactive network, which is then partitioned into topic-specific communities using its topological structure. Within each community, dense subgraphs are mined on the basis of time dimension, capturing temporal dynamics.
@@
However, current dynamic social network community discovery algorithms still have many limitations: (a) The explosive growth of social network data increases computational complexity, memory usage, and runtime. Many algorithms require recalculating at each time step, causing the time complexity to rise significantly with the number of steps, hindering real-time performance. (b) Some algorithms neglect temporal correlations and fail to capture the true evolution of communities and complex structures such as nesting, overlap, and hierarchy, leading to the loss of critical information. (c) Algorithms often struggle with the dynamic nature of social networks, exhibiting poor adaptability to structural changes and sensitivity to parameter adjustments. In addition, they are vulnerable to noise and abnormal data, compromising accuracy measured in purity and NMI and reliability. Considering these limitations, to fully extract the dynamic features of the network and the interaction information between nodes, and to deeply mine the dynamic features of the network, we propose a community discovery algorithm based on spatiotemporal graph embedding (GTENN). This algorithm employs Graph Convolutional Neural Networks (GCNs) to aggregate latent spatial information, effectively representing the embedding of nodes in space. The temporal evolutions of nodes are then modeled using Gated Recurrent Units (GRU), thereby solving problems such as node dynamism and relationship transmission. Finally, a Self-Organizing Map (SOM) is applied to cluster dynamic network representations and identify community affiliations of nodes.
@@ 3.1 Definitions
Definition 1 (Dynamic Network).
Let $V=\{v_{1},v_{2},\dots,v_{n}\}$ be the set of nodes and $E=\{e_{ij}\mid i\neq j,i,j\in\{1,\dots,n\}\}$ the set of potential edges in a dynamic network. At each time step $t$, the snapshot network is defined as $g_{t}=(V,E_{t})$, where $E_{t}\subseteq E$ varies over time while the node set $V$ remains fixed.
The full dynamic network is then represented as a sequence of time-varying snapshots:
$G=\{g_{1},g_{2},\dots,g_{T}\},\quad 1\leq t\leq T.$
Each snapshot $g_{t}$ is associated with an adjacency matrix $A_{t}\in\mathbb{R}^{n\times n}$, where $A_{t}(i,j)$ encodes the edge weight from node $v_{i}$ to node $v_{j}$ at time $t$. Unless otherwise specified, all graphs are assumed to be undirected, i.e., $A_{t}=A_{t}^{\top}$, and unweighted, i.e., $A_{t}(i,j)\in\{0,1\}$.
Definition 2 (Dynamic Community).
We use $C_{t}=\{c_{t}^{1},c_{t}^{2},\dots,c_{t}^{p}\mid\forall t\in\{1,\dots,T\}\}$ to denote the $p$ communities at time slice network $g_{t}$, where each $c_{t}^{i}\subseteq V$ is a subset of nodes belonging to the $i$-th community at time $t$.
The dynamic community structure over the entire dynamic network $G$ is represented as:
$C=\{C_{1},C_{2},\dots,C_{T}\}.$
Definition 3 (Graph Embedding).
Graph embedding is a mapping function $\Gamma$ that projects each node $v_{i}$ at time $t$ into a low-dimensional latent space:
$\Gamma:v_{i}^{t}\mapsto f_{i}^{t}\in\mathbb{R}^{d},\quad 1\leq i\leq n,\;1\leq t\leq T,\;d\ll n.$
Here, $f_{i}^{t}$ is the embedding vector for node $v_{i}$ at time $t$, and $d$ is the embedding dimension. The matrix of all node embeddings at time $t$ is denoted by $F_{t}=[f_{1}^{t},f_{2}^{t},\dots,f_{n}^{t}]^{\top}\in\mathbb{R}^{n\times d}$, and the full sequence of embedding matrices is denoted as:
$F_{S}=\{F_{1},F_{2},\dots,F_{T}\}.$
Definition 4 (Adjacency Matrix).
Given a dynamic network $G=\{g_{t}\}_{t=1}^{T}$, each snapshot $g_{t}=(V,E_{t})$ consists of a fixed set of nodes $V$ and a set of edges $E_{t}$ at time $t$. The adjacency matrix at time $t$ is defined as
$A_{t}(i,j)=\begin{cases}1,&\text{if }(v_{i},v_{j})\in E_{t},\\ 0,&\text{otherwise},\end{cases}\quad\forall v_{i},v_{j}\in V.$
Since we consider undirected graphs [^1] Self-loops are not included unless explicitly added for normalization purposes., the adjacency matrix is symmetric, i.e., $A_{t}(i,j)=A_{t}(j,i)$.
@@
$G$ Dynamic network (sequence of graphs)
$g_{t}$ Graph snapshot at time $t$
$A_{t}$ Adjacency matrix of $g_{t}$
$\hat{A}_{t}$ Normalized adjacency matrix at time $t$
$t$ Current time step
$T$ Total number of time steps
$n$ Number of nodes
@@
$d$ Dimension of embedding vector
$C$ Dynamic community collection $\{C_{t}\}_{t=1}^{T}$
$F_{t}$ Embedding matrix at time $t\in\mathbb{R}^{n\times d}$
$F_{S}$ Sequence of all embeddings over time
$f_{i}^{t}$ Embedding vector of node $v_{i}$ at time $t$
$\Gamma$ Embedding function
Table 1: Notations used throughout the paper, including dynamic network representations, time-indexed graph structures, embedding dimensions, and community-related variables.
3.2 The GTENN Algorithm
The dynamic network community discovery process based on spatiotemporal graph embedding, shown in Fig. 1, consists of three main components: spatial feature extraction, temporal feature extraction, and community discovery. In order to fully utilize the spatial information of the nodes in the network, we first employ GCNs to learn the feature representation of the nodes from a spatial perspective and aggregate the social information of users who communicate with each other. Next, we share the GCNs’ weight parameters using GRU to maintain the dynamics of spatial nodes and extract the features of dynamic nodes, addressing node dynamics and long-term dependencies to form the feature representation of the dynamic network. Finally, the SOM algorithm is applied to analyze the network embedding and perform community discovery for dynamic artificial and real social networks. We dub this process as the GTENN algorithm.
Figure 1: Framework of the proposed GTENN algorithm. The process integrates three main components: (i) spatial feature extraction using GCNs to learn node representations from network topology, (ii) temporal feature extraction via GRU to capture node dynamics and long-term dependencies by sharing GCN weights across time, and (iii) community discovery using SOM to analyze the learned embeddings and identify evolving communities in dynamic networks.
@@ 3.3 Spatial Feature Extraction
In the spatial feature extraction stage, we start with node interaction learning and use GCNs to learn the topological structure information of nodes in the snapshot network and obtain the embeddings of nodes. This process involves two key aspects: space embedding learning and weight parameter sharing.
@@ 3.3.1 Space Embedding Learning
GCNs consist of multiple layers of graph convolutions that aggregate neighborhood features. At time step $t$, in the $l$-th GCN layer, the model takes as input the node representation matrix $F_{t}^{l-1}\in\mathbb{R}^{n\times d}$, the snapshot adjacency matrix $A_{t}\in\mathbb{R}^{n\times n}$, and a learnable weight matrix $W_{t}^{l-1}\in\mathbb{R}^{d\times d^{\prime}}$. The updated node representation matrix is computed as:
$F_{t}^{l}=\Gamma\left(F_{t}^{l-1},A_{t},W_{t}^{l-1}\right)=\sigma\left(\hat{A}_{t}F_{t}^{l-1}W_{t}^{l-1}\right),$ (1)
where $\sigma(\cdot)$ is the activation function, and in this paper we adopt the LeakyReLU function. The matrix $\hat{A}_{t}\in\mathbb{R}^{n\times n}$ is the symmetrically normalized adjacency matrix used as a graph convolution filter, computed as follows.
We first augment the original adjacency matrix $A_{t}$ with self-loops by adding the identity matrix:
$\tilde{A}_{t}=A_{t}+\mathbb{I},$ (2)
where $\mathbb{I}\in\mathbb{R}^{n\times n}$ is the identity matrix. Then, we compute the degree matrix $\tilde{D}_{t}\in\mathbb{R}^{n\times n}$ of $\tilde{A}_{t}$ as:
$\tilde{D}_{t}=\operatorname{diag}\left(\sum_{j=1}^{n}\tilde{A}_{t}(i,j)\right).$ (3)
The normalized adjacency matrix $\hat{A}_{t}$ is then given by:
$\hat{A}_{t}=\tilde{D}_{t}^{-1/2}\tilde{A}_{t}\tilde{D}_{t}^{-1/2}.$ (4)
This normalization ensures that neighborhood feature aggregation is appropriately scaled and robust to variations in node degree. The resulting matrix $\hat{A}_{t}$ acts as a localized spectral filter, approximating graph convolution via eigenspace regularization in a computationally efficient manner.
At time $t$, the input to the first layer is a randomly initialized node feature matrix $F_{t}^{0}\in\mathbb{R}^{n\times d}$, where each row $f_{t}^{i}\in\mathbb{R}^{d}$ represents the latent embedding of node $v_{i}$. This initialization is adopted due to the lack of observed node attributes in the dynamic network.
The output of the final GCN layer at time $t$, denoted as $F_{t}^{l}$, encodes both local topology and structural semantics within the snapshot $g_{t}$. Across time, the set of snapshot-specific node embeddings is represented as:
$F_{t}=\left\{f_{t}^{1},f_{t}^{2},\ldots,f_{t}^{n}\right\},\quad 1\leq t\leq T.$ (5)
@@ 3.3.2 Weight Parameter Sharing
Given the dynamic nature of the network, dynamic update strategies are incorporated into the graph convolutional layers. This is necessary because any change in nodes or edges triggers an adjustment of the weight parameters in the convolutional layers to adapt to the evolving network structure. Therefore, we follow the method of EvolveGCNs[15] and use GRU to update the weight parameters of GCNs at the previous time step, which serves as the weight of GCNs at the current time step, to adapt to the development of dynamic networks. The weight update strategy for the $l$-th layer at time t is as follows:
@@
The GCNs module aggregates information between nodes and their neighbors, while the GRU adaptively adjusts weight parameters over time. Together, these modules dynamically capture the evolving neighborhood information of the nodes. Thus, operations in the spatial feature extraction stage can be formally represented as
@@
Figure 2: Spatial feature extraction stage of GTENN, where GCNs learn node features by aggregating structural information.